# RobustMerge: Parameter-Efficient Model Merging for MLLMs with Direction Robustness

### v2

@@ Abstract
Fine-tuning pre-trained models with custom data leads to numerous expert models on specific tasks. Merging models into one universal model to empower multi-task ability refraining from data leakage has gained popularity. With the expansion in data and model size, parameter-efficient tuning becomes the common practice for obtaining task-specific models efficiently. However, few methods are dedicated to efficient merging, and existing methods designed for full fine-tuning merging fail under efficient merging. To address the issue, we analyze from low-rank decomposition and reveal that direction robustness during merging is crucial for merging efficient modules. We furthermore uncover that compensating for the gap between stark singular values contributes to direction robustness. Therefore, we propose RobustMerge, a training-free parameter-efficient merging method with complementary parameter adaptation to maintain direction robustness. Specifically, we (1) prune parameters and scale coefficients from inter-parameter relation for singular values to maintain direction stability away from task interference, and (2) perform cross-task normalization to enhance unseen task generalization. We establish a benchmark consisting of diverse multimodal tasks, on which we conduct experiments to certify the outstanding performance and generalizability of our method. Additional studies and extensive analyses further showcase the effectiveness. Code is available at https://github.com/AuroraZengfh/RobustMerge.
@@ 1 Introduction
Despite its popularity, crucial problems for model merging remain unsolved, restricting its real-world deployment. First, with larger model sizes like multimodal large language models (MLLMs) and massive data, parameter-efficient tuning (PEFT) hu2022lora has become the most popular and effective tuning approach for large models. However, existing model merging methods focus on full fine-tuning (FFT) techniques yadav2024ties ; guodong2024pcb , which struggle in distribution shift and undergo performance drops when directly applied to parameter-efficient model merging, as is illustrated in Fig. 2. Moreover, another issue lies in that current high-performance methods rely on extra information of seen tasks (e.g., validation data yang2024adamerging , extra storage huang2024emr ) to boost the performance. Therefore, they can only handle seen tasks and fail to generalize to unseen tasks, questioning their robustness and extensibility in real-world scenarios, as is concluded in Tab. 2. The most related work is LoraHub huang2023lorahub . However, its requirement for coefficient optimization through test time adaptation severely hinders its application.
@@
Inspired by the stated shortcomings, we aim to develop a merging algorithm for parameter-efficient modules with generalizability. First, we analyze the reason behind the performance drop. We observe (1) stark singular values and (2) a distinct wider distribution in efficient parameters that differ from full fine-tuning. Moreover, starting from the perspective of low-rank decomposition, we reveal that direction robustness, i.e., maintaining directions of singular values, is crucial for efficient merging. From the above observations, we propose RobustMerge, a novel parameter-efficient method for high-performance merging of multimodal large models and introduce effective complementary [^1] In contrast to individual, we use the term to distinguish between subspace multiplication (along $r$ dimension) and original multiplication (along $d_{i}/d_{o}$ dimension) of matrix. parameter adaptation to maintain directions for performance enhancement. Concretely, we prune ineffective parameters and construct scaling coefficients from inter-parameter relation directly on LoRA components to mitigate interference between tasks aroused from stark singular values difference. Additionally, we perform cross-task normalization to balance tasks of different data scales and enhance unseen task generalization. It is notable that our method is free from any additional data or storage and does not require explicit decomposition, which equips the method with more flexibility and efficiency.
@@ 2 Related Work
Multimodal large language models. With the surge in data volume and model size, large language models (LLMs) radford2019language ; touvron2023llama ; alayrac2022flamingo have shown their powerful performance. They are constructed with decoder-only blocks and respond to inputs in an auto-regressive way, which shows their potential on both classification wang2018glue and generative tasks goyal2017vqa . Furthermore, multimodal large language models (MLLMs) enhance large models with vision perception ability. They obtain visual features with a vision encoder and align image-text features with a cross-modality module liu2024visual ; li2023blip and so on. Current research on large models is dedicated to directly fine-tuning one independent model with task-specific data to get better results. Rather than improving the performance of a certain domain, we focus on integrating models into one model to boost efficiency and handle multiple tasks simultaneously.
@@
Parameter-efficient tuning. When fine-tuning a pre-trained model with task-specific data, training the whole model would not only disrupt the representations obtained from billions of data but also become resource-intensive. To address the issue, parameter-efficient tuning han2024parameter is introduced to refrain from fine-tuning the whole model. It typically trains lightweight modules to make the model adapt to downstream tasks and achieves competitive results compared to full fine-tuning models. Various efficient tuning techniques have been explored, like prompt learning jia2022visual ; khattak2023maple ; zeng2025local , adapter learning including LoRA hu2022lora ; wu2024moslora ; liu2024dora ; meng2024pissa , (IA)3 liu2022few and so on. In this paper, we focus on LoRA, as it is the most commonly utilized PEFT method and has demonstrated its usefulness in various fields zhu2025trustlora ; guo2025federated especially for large models liu2024visual .
@@
Model merging. Model merging yang2024model ; sung2023empirical refers to merging multiple models of different capabilities to handle multi-task learning with one universal model jin2023regmean ; matena2022merging . Task Arithmetic ilharco2023editing presents a paradigm that obtains task vectors from subtracting a pre-trained model from fine-tuned models and treats model merging as arithmetic operations of task vectors. It has gained widespread attention in various fields tang2024wemoe . Ties-merging yadav2024ties trims and elects signs to reduce interference. DARE yu2024dare randomly drops parameters and rescales the remaining ones to approximate the original embedding. PCB-merging guodong2024pcb introduces parameter adjustment with competition balancing to address potential conflicts. However, most of them focus on merging models with FFT on classification tasks deng2009imagenet , and the distribution shift prevents their ability to acquire satisfying performance tang2024parameter . Some recent works liu2024pruning ; liu2024checkpoint also focus on merging checkpoints during pre-training to enhance downstream performance. By contrast, we concentrate on parameter-efficient merging with multimodal tasks.
@@ 3.1 Preliminary and Notations
Model merging targets at combining multiple models of the same structure $\{\theta_{1},\cdots,\theta_{N}\}$ that are fine-tuned from pre-trained model $\theta_{pre}$ into one new model $\theta_{m}$ and maintaining multi-task ability in a training-free manner. Existing full fine-tuning (FFT) methods follow the paradigm proposed by Task Arithmetic (TA) ilharco2023editing . Typically, they construct the task vector by performing a subtraction operation $\tau_{n}=\theta_{n}-\theta_{pre}\in\mathbb{R}^{d}$, conduct merging algorithm on task vector subspace and obtain the final merged model by adding pre-trained model, i.e., $\theta_{m}=\theta_{pre}+\lambda\sum_{n=1}^{N}\Phi(\tau_{n})$, where $\Phi(\cdot)$ stands for merging algorithm.
Parameter-efficient model merging differs from traditional model merging, as the backbone of the foundation model is frozen and the updated matrices to be merged are randomly initialized. Consequently, we use $\Delta\boldsymbol{\mathrm{W}}$ to represent merging modules for simplification, and exploit the model merging method on these parameter-efficient modules, i.e., $\boldsymbol{\mathrm{W}}_{m}=\boldsymbol{\mathrm{W}_{0}}+\lambda\sum_{n=1}^{N}\Phi(\Delta\boldsymbol{\mathrm{W}}_{n})$.
@@ 3.3 RobustMerge: Parameter-Efficient Model Merging for MLLMs
 $\Delta\widetilde{\boldsymbol{\mathrm{W}}}_{n}\leftarrow\widetilde{\boldsymbol{\mathrm{B}}}_{n}\cdot\widetilde{\boldsymbol{\mathrm{S}}}_{n}\cdot\widetilde{\boldsymbol{\mathrm{A}}}_{n}$
@@
  $\boldsymbol{\mathrm{W}}\leftarrow\boldsymbol{\mathrm{W}_{0}}+\lambda\sum_{n=1}^{N}\Delta\widetilde{\boldsymbol{\mathrm{W}}}_{n}$
@@
$\Delta\widetilde{\boldsymbol{\mathrm{W}}}_{n}=\widetilde{\boldsymbol{\mathrm{B}}}_{n}\cdot\widetilde{\boldsymbol{\mathrm{S}}}_{n}\cdot\widetilde{\boldsymbol{\mathrm{A}}}_{n},n=1,\cdots,N,$ (5)
@@ 4.3 Ablation Study and Further Analysis
Influence of pruning rate. As is revealed in Sec. 3.2, parameters of small values have less influence on the direction robustness of low-rank decomposition. Therefore, we prune parameters according to the magnitude to facilitate the merging procedure. To further validate this point of view, we gradually increase the pruning rate and show the variation of average performance in Fig. 7a. It is elucidated that: (1) With the pruning rate increases, the performance gradually boosts due to reduced interference between tasks; (2) Finally, the performance undergoes a sharp decrease as pruning larger parameters significantly influences directions and task knowledge. Consequently, the results are in accordance with the aforementioned analysis, underlining the utility of the ineffective parameter pruning strategy in our method.
Figure 9: Influence of complementary parameter scaling. Coefficients solely dependent on specific module ($\boldsymbol{\mathrm{A}}$, $\boldsymbol{\mathrm{B}}$ or both) perform inferior to adaptive coefficients with inter-parameter relation.
@@
Figure 10: Similarity of singular vector and ratio of singular value on different models.
Further analysis of direction robustness. We formulate the evaluation metrics for measuring robustness of model merging to better understand of the effectiveness of merging methods. We use the average similarity of each corresponding singular vector between task-specific models and merged models as the criterion, which reflects the direction deviation in merging. Larger similarity means the direction possesses more robustness and is not prone to change direction during merging, thereby maintaining performance of specific task. Moreover, we also incorporate the ratio of singular value between merged and original models to comprehensively reflect the degree of specific knowledge learning. As is depicted in Fig. 4, the largest singular value displays more direction robustness, so we divide the value into the largest and the average of the remaining parts in Fig. 10 for better illustration.
@@ 5 Conclusion
Limitations and future work. We do not validate the proposed method on different structures and more types of tasks. However, since our method is a model-agnostic and task-agnostic post-processing merging algorithm, we believe this will not be a bottleneck given the large number of existing models on platforms like Huggingface. Also, we propose the concept of direction robustness in parameter-efficient merging, but we do not design a specific algorithm on decomposed matrices for the purposes of efficiency. We believe they would be promising directions that are left for future development.
@@ A.1 Background of Singular Value Decomposition
Denote the parameter-efficient module $\boldsymbol{\mathrm{W}}=\boldsymbol{\mathrm{B}}\times\boldsymbol{\mathrm{A}}$, $\boldsymbol{\mathrm{W}}\in\mathbb{R}^{n\times n}$ and $\boldsymbol{\mathrm{W}}$ is a low-rank matrix, i.e., Rank($\boldsymbol{\mathrm{W}}$) = $r$, $r\ll n$. Singular values of matrix $\boldsymbol{\mathrm{W}}$ are $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}>0$.
@@
$=$ $\sigma_{1}u_{1}v_{1}^{T}+\sigma_{2}u_{2}v_{2}^{T}+\cdots+\sigma_{r}u_{r}v_{r}^{T},$
@@ A.2 Theoretical Analysis in Merging
where $\sigma_{ij},u_{ij}$ and $v_{ij}$ represents the $i^{th}$ singular value/left singular vector/right singular vector of the $j^{th}$ low-rank matrix, respectively.
Given the practical significance of singular vectors in LoRA, consider two permutations of 1 to $r$, i.e., $(\overline{1}),(\overline{2}),\cdots,(\overline{r}),$ and $(\underline{1}),(\underline{2}),\cdots,(\underline{r}),$, merging the two modules can therefore be rewritten as:
@@
where each pairwise subscript $\{(\overline{i}),(\underline{i})\},i=1,\cdots,r$ stand for similar singular components, i.e., similar knowledge of two different matrices in low-rank space.
Empirically, $\boldsymbol{U}$ contains more general knowledge in low-rank space with larger similarity across tasks. Consequently, the merging process can mathematically be expressed as merging each of task-specific vectors in low-rank space:
@@
It can be seen from the above derivation that due to the difference in original direction and magnitude, the singular values with larger magnitude are more likely to determine the direction and magnitude of the merged singular vector. As a result, the change in singular vector angle will vary from the perspective of singular value vectors belonging to different tasks due to the stark difference in singular values, e.g., for task A, the direction angle change for vector 1 is small while angle change for vector 2 is relatively large, and the situation is just the opposite for task B. Therefore, the key for merging would be maintaining direction robustness for vectors with small singular values. Without the loss of generality, the derivation can be extended to merging any number of models.

### v5

@@ Abstract
Fine-tuning pre-trained models with custom data leads to numerous expert models on specific tasks. Merging models into one universal model to empower multi-task ability refraining from data leakage has gained popularity. With the expansion in data and model size, parameter-efficient tuning becomes the common practice for obtaining task-specific models efficiently. However, few methods are dedicated to efficient merging, and existing methods designed for full fine-tuning merging fail under efficient merging. To address the issue, we analyze from low-rank decomposition and reveal that direction robustness during merging is crucial for merging efficient modules. We furthermore uncover that compensating for the gap between stark singular values contributes to direction robustness. Therefore, we propose RobustMerge, a training-free parameter-efficient merging method with complementary parameter adaptation to maintain direction robustness. Specifically, we (1) prune parameters and scale coefficients from inter-parameter relations for singular values to maintain direction stability away from task interference, and (2) perform cross-task normalization to enhance unseen task generalization. We establish a benchmark consisting of diverse multimodal tasks, on which we conduct experiments to certify the outstanding performance and generalizability of our method. Additional studies and extensive analyses further showcase the effectiveness. Code is available at https://github.com/AuroraZengfh/RobustMerge.
@@ 1 Introduction
Despite its popularity, crucial problems for model merging remain unsolved, restricting its real-world deployment. First, with larger model sizes like multimodal large language models (MLLMs) and massive data, parameter-efficient tuning (PEFT) hu2022lora has become the most popular and effective tuning approach for large models. However, existing model merging methods focus on full fine-tuning (FFT) techniques yadav2024ties ; guodong2024pcb , which struggle with distribution shift and undergo performance drops when directly applied to parameter-efficient model merging, as is illustrated in Fig. 2. Moreover, another issue lies in that current high-performance methods rely on extra information of seen tasks (e.g., validation data yang2024adamerging , extra storage huang2024emr ) to boost the performance. Therefore, they can only handle seen tasks and fail to generalize to unseen tasks, questioning their robustness and extensibility in real-world scenarios, as is concluded in Tab. 2. The most related work is LoraHub huang2023lorahub . However, its requirement for coefficient optimization through test-time adaptation severely hinders its application.
@@
Inspired by the stated shortcomings, we aim to develop a merging algorithm for parameter-efficient modules with generalizability. First, we analyze the reason behind the performance drop. We observe (1) stark singular values and (2) a distinct wider distribution in efficient parameters that differ from full fine-tuning. Moreover, starting from the perspective of low-rank decomposition, we reveal that direction robustness, i.e., maintaining directions of singular values, is crucial for efficient merging. From the above observations, we propose RobustMerge, a novel parameter-efficient method for high-performance merging of multimodal large models and introduce effective complementary [^1] In contrast to individual, we use the term to distinguish between subspace multiplication (along $r$ dimension) and original multiplication (along $d_{i}/d_{o}$ dimension) of matrix. parameter adaptation to maintain directions for performance enhancement. Concretely, we prune ineffective parameters and construct scaling coefficients from inter-parameter relations directly on LoRA components to mitigate interference between tasks aroused from stark singular values difference. Additionally, we perform cross-task normalization to balance tasks of different data scales and enhance unseen task generalization. It is notable that our method is free from any additional data or storage and does not require explicit decomposition, which equips the method with more flexibility and efficiency.
@@ 2 Related Work
Multimodal large language models. With the surge in data volume and model size, large language models (LLMs) radford2019language ; touvron2023llama ; alayrac2022flamingo have shown their powerful performance. They are constructed with decoder-only blocks and respond to inputs in an auto-regressive way, which shows their potential in both classification wang2018glue and generative tasks goyal2017vqa . Furthermore, multimodal large language models (MLLMs) enhance large models with vision perception ability. They obtain visual features with a vision encoder and align image-text features with a cross-modality module liu2024visual ; li2023blip and so on. Current research on large models is dedicated to directly fine-tuning one independent model with task-specific data to get better results. Rather than improving the performance of a certain domain, we focus on integrating models into one model to boost efficiency and handle multiple tasks simultaneously.
@@
Parameter-efficient tuning. When fine-tuning a pre-trained model with task-specific data, training the whole model would not only disrupt the representations obtained from billions of data but also become resource-intensive. To address the issue, parameter-efficient tuning han2024parameter is introduced to refrain from fine-tuning the whole model. It typically trains lightweight modules to make the model adapt to downstream tasks and achieves competitive results compared to full fine-tuning models. Various efficient tuning techniques have been explored, like prompt learning jia2022visual ; khattak2023maple ; zeng2025local , adapter learning including LoRA hu2022lora ; wu2024moslora ; liu2024dora ; meng2024pissa ; liu2025rora , (IA)3 liu2022few and so on. In this paper, we focus on LoRA, as it is the most commonly utilized PEFT method and has demonstrated its usefulness in various fields zhu2025trustlora ; guo2025federated especially for large models liu2024visual .
@@
Model merging. Model merging yang2024model ; sung2023empirical ; li2024map refers to merging multiple models of different capabilities to handle multi-task learning with one universal model jin2023regmean ; matena2022merging . Task Arithmetic ilharco2023editing presents a paradigm that obtains task vectors from subtracting a pre-trained model from fine-tuned models and treats model merging as arithmetic operations of task vectors. It has gained widespread attention in various fields tang2024wemoe . Ties-merging yadav2024ties trims and elects signs to reduce interference. DARE yu2024dare randomly drops parameters and rescales the remaining ones to approximate the original embedding. PCB-merging guodong2024pcb introduces parameter adjustment with competition balancing to address potential conflicts. However, most of them focus on merging models with FFT on classification tasks deng2009imagenet , and the distribution shift prevents their ability to acquire satisfying performance tang2024parameter . Some recent works liu2024pruning ; liu2024checkpoint also focus on merging checkpoints during pre-training to enhance downstream performance. By contrast, we concentrate on parameter-efficient merging with multimodal tasks.
@@ 3.1 Preliminary and Notations
Model merging targets at combining multiple models of the same structure $\{\theta_{1},\cdots,\theta_{N}\}$ that are fine-tuned from pre-trained model $\theta_{pre}$ into one new model $\theta_{m}$ and maintaining multi-task ability in a training-free manner. Existing full fine-tuning (FFT) methods follow the paradigm proposed by Task Arithmetic (TA) ilharco2023editing . Typically, they construct the task vector by performing a subtraction operation $\tau_{n}=\theta_{n}-\theta_{pre}\in\mathbb{R}^{d}$, conduct a merging algorithm on the task vector subspace, and obtain the final merged model by adding the pre-trained model, i.e., $\theta_{m}=\theta_{pre}+\lambda\sum_{n=1}^{N}\Phi(\tau_{n})$, where $\Phi(\cdot)$ stands for the merging algorithm.
Parameter-efficient model merging differs from the traditional model merging, as the backbone of the foundation model is frozen and the updated matrices to be merged are randomly initialized. Consequently, we use $\Delta\boldsymbol{\mathrm{W}}$ to represent merging modules for simplification, and exploit the model merging method on these parameter-efficient modules, i.e., $\boldsymbol{\mathrm{W}}_{m}=\boldsymbol{\mathrm{W}_{0}}+\lambda\sum_{n=1}^{N}\Phi(\Delta\boldsymbol{\mathrm{W}}_{n})$.
@@ 3.3 RobustMerge: Parameter-Efficient Model Merging for MLLMs
$\Delta\widetilde{\boldsymbol{\mathrm{W}}}\leftarrow\lambda\ \sum_{i=1}^{N}(\widetilde{\boldsymbol{\mathrm{B}}}_{n}\cdot\widetilde{\boldsymbol{\mathrm{S}}}_{n})\cdot\sum_{i=1}^{N}\widetilde{\boldsymbol{\mathrm{A}}}_{n}$
  $\boldsymbol{\mathrm{W}}\leftarrow\boldsymbol{\mathrm{W}_{0}}+\Delta\widetilde{\boldsymbol{\mathrm{W}}}$
@@
$\Delta\widetilde{\boldsymbol{\mathrm{W}}}=\lambda\ {\textstyle\sum_{i=1}^{N}}(\widetilde{\boldsymbol{\mathrm{B}}}_{n}\cdot\widetilde{\boldsymbol{\mathrm{S}}}_{n})\cdot{\textstyle\sum_{i=1}^{N}}\widetilde{\boldsymbol{\mathrm{A}}}_{n},$ (5)
@@ 4.3 Ablation Study and Further Analysis
Influence of pruning rate. As is revealed in Sec. 3.2, parameters of small values have less influence on the direction robustness of low-rank decomposition. Therefore, we prune parameters according to the magnitude to facilitate the merging procedure. To further validate this point of view, we gradually increase the pruning rate and show the variation of average performance in Fig. 7a. It is elucidated that: (1) As the pruning rate increases, the performance gradually boosts due to reduced interference between tasks; (2) Finally, the performance undergoes a sharp decrease as pruning larger parameters significantly influences directions and task knowledge. Consequently, the results are in accordance with the aforementioned analysis, underlining the utility of the ineffective parameter pruning strategy in our method.
Figure 9: Influence of complementary parameter scaling. Coefficients solely dependent on specific module ($\boldsymbol{\mathrm{A}}$, $\boldsymbol{\mathrm{B}}$ or both) perform inferior to adaptive coefficients with inter-parameter relations.
@@
Figure 10: Similarity of singular vector and ratio of singular value on different merging techniques.
Further analysis of direction robustness. We formulate the evaluation metrics for measuring direction robustness to better understand parameter-efficient model merging. We use the average similarity of each corresponding singular vector between task-specific models and merged models as the criterion, which reflects the direction deviation in merging. Larger similarity means the direction possesses more robustness and is not prone to changing direction during merging, thereby maintaining the performance of a specific task. Moreover, we also incorporate the ratio of singular value between merged and original models to comprehensively reflect the degree of specific knowledge learning. As is depicted in Fig. 4, the largest singular value displays more direction robustness, so we divide the value into the largest and the average of the remaining parts in Fig. 10 for better illustration of different model merging approaches.
@@ 5 Conclusion
Limitations and future work. We do not validate the method on more structures and tasks. However, since our method is a model-agnostic and task-agnostic post-processing algorithm, this will not be a bottleneck given numerous models on platforms like Huggingface. Also, we propose the concept of direction robustness in parameter-efficient merging, but we do not design a specific algorithm on decomposed matrices for the purposes of efficiency. We believe they would be promising directions that are left for future development in various downstream areas of parameter-efficient learning.
@@ A.1 Background of Singular Value Decomposition
Denote the parameter-efficient module $\boldsymbol{\mathrm{W}}=\boldsymbol{\mathrm{B}}\times\boldsymbol{\mathrm{A}}$, $\boldsymbol{\mathrm{W}}\in\mathbb{R}^{n\times n}$, and $\boldsymbol{\mathrm{W}}$ as a low-rank matrix, i.e., Rank($\boldsymbol{\mathrm{W}}$) = $r$, $r\ll n$. The singular values of the matrix $\boldsymbol{\mathrm{W}}$ are $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}>0$.
@@
$=$ $\sigma_{1}u_{1}v_{1}^{T}+\sigma_{2}u_{2}v_{2}^{T}+\cdots+\sigma_{r}u_{r}v_{r}^{T}.$
@@ A.2 Theoretical Analysis in Merging
where $\sigma_{ij},u_{ij}$ and $v_{ij}$ represent the $i^{th}$ singular value/left singular vector/right singular vector of the $j^{th}$ low-rank matrix, respectively.
Given the practical significance of singular vectors in LoRA, consider two permutations of 1 to $r$, i.e., $(\overline{1}),(\overline{2}),\cdots,(\overline{r}),$ and $(\underline{1}),(\underline{2}),\cdots,(\underline{r})$, merging the two modules can therefore be rewritten as:
@@
where each pairwise subscript $\{(\overline{i}),(\underline{i})\},i=1,\cdots,r$ stands for similar singular components, i.e., similar knowledge of two different matrices in low-rank space.
Empirically, $\boldsymbol{U}$ contains more general knowledge in low-rank space with larger similarity across tasks. Consequently, the merging process can mathematically be expressed as merging each of the task-specific vectors in low-rank space:
@@
It can be seen from the above derivation that due to the difference in original direction and magnitude, the singular values with larger magnitude are more likely to determine the direction and magnitude of the merged singular vector. As a result, the change in singular vector angle will vary from the perspective of singular value vectors belonging to different tasks due to the stark difference in singular values, e.g., for task A, the direction angle change for vector 1 is small while the angle change for vector 2 is relatively large, and the situation is just the opposite for task B. Therefore, the key for merging would be maintaining direction robustness for vectors with small singular values. Without the loss of generality, the derivation can be extended to merging any number of models.